OpenStudy (anonymous):

Help please! If x varies inversely with y and x = 6 when y = 10, find y when x = 8.

OpenStudy (anonymous):

12

OpenStudy (anonymous):

10

OpenStudy (anonymous):

That's not an option, @kidsosa23 These are the choices y=4/15 y=4.8 y=7.5 y=40/3

OpenStudy (anonymous):

7.5

OpenStudy (anonymous):

40/3 is 13.3

OpenStudy (anonymous):

y=7.5 is the answer

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

yep can i have my medal

OpenStudy (anonymous):

np

OpenStudy (anonymous):

By all means

OpenStudy (anonymous):

Does anyone have time for a few more?

OpenStudy (anonymous):

i do

OpenStudy (anonymous):

so do i

OpenStudy (anonymous):

If I varies inversely with R and I = 0.75 when R = 14, which equation should be used to show this relationship? I=R/10.5 I=10.5/R I=10.5*R IR=1.05

OpenStudy (anonymous):

@chapmcom @blake177

OpenStudy (anonymous):

IR=1.05

OpenStudy (anonymous):

Thank you, again. Here's the next one, xy = –25. What is the constant of variation for this relationship? 1/4 -0.4 -2.5 -25

OpenStudy (anonymous):

-25

OpenStudy (anonymous):

Much gratitude. This will be the last one. :) V varies inversely with T and V = 12 when T = 6. Which equation shows this relationship? V=72/T V=2T V=2/T V=72T

OpenStudy (anonymous):

V=2T

OpenStudy (anonymous):

Thank you again for all your help. Please feel free to contact me when you need assistance. :)

OpenStudy (anonymous):

no problem glad to help

OpenStudy (anonymous):

So it turns out that I got a less than passing grade on the test, but don't worry. It was more my fault than yours. I will post all the correct answers now.

OpenStudy (anonymous):

thats no right all of those were right

OpenStudy (anonymous):

i double checked em on my calculator

OpenStudy (anonymous):

Y=7/5 I=10.5/R -25 V=72/T

OpenStudy (anonymous):

That's alright. It's the thought that counts.

OpenStudy (anonymous):

well sorry

OpenStudy (anonymous):

Don't stress. If I had done it by myself, I would have gotten a zero, so..

OpenStudy (anonymous):

btw love your picture lol

OpenStudy (anonymous):

Haha, Thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

well i gotta go